OpenStudy (anonymous):

how to find the derivative of (4x+1)(1-x)^3

OpenStudy (freckles):

Product rule would need to be applied.

OpenStudy (freckles):

Then Chain rule and some other rules.

OpenStudy (freckles):

but yeah the first one to begin with is product

OpenStudy (freckles):

\[\frac{d}{dx}(4x+1)(1-x)^3 \\ = (1-x)^3 \frac{d}{dx}(4x+1)+ (4x+1) \frac{d}{dx}(1-x)^3\] see if you can continue

OpenStudy (anonymous):

\[(4 x+1) (1-x)^3=-4 x^4+11 x^3-9 x^2+x+1 \]Calculate the derivative term by term.\[-16 x^3+33 x^2-18 x+1 \]